The quality factor of a weakly damped
oscillator of frequency 200 Hz is 1000.
Calculate the time in which its energy
becomes 20% of the initial value.
1
Expert's answer
2019-05-27T10:53:22-0400

Given:

"Q = 1000"

"f = 200 Hz"

"\frac{E_t}{E_0} = 20\% = 0.2"

Solution:

The quality factor for weakly damped oscillator is :


"Q = \frac{\omega _0}{\gamma}"

So, damping rate is:


"\gamma = \frac{\omega _ 0 }{Q}"

Energy of such oscillator is:


"E(t) = E_0 e^{-\gamma t}"

Solving for t:


"E_t = E_0 e^{-\frac{\omega _0 t}{Q}}"

"\frac{E_t}{E_0} = e^{-\frac{\omega _0 t}{Q}}"

"ln(\frac{E_t}{E_0} )= -\frac{\omega _0 t}{Q}"

Finally:


"t = -\frac{Q}{\omega _0} ln(\frac{E_t}{E_0}) = -\frac{Q}{2 \pi f} ln(\frac{E_t}{E_0})"

Calculating:


"t \approx -\frac{1000}{2 * 3.14 * 200} ln(0.2) \approx 1.6 *\frac{1000}{1256} = 1.27 sec"

Answer: 1.27 sec